\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
% \MRhref is called by the amsart/book/proc definition of \MR.
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}
\begin{thebibliography}{Her11b}

\bibitem[BL94]{B-L94}
Laure Barthel and Ron Livn{\'e}, \emph{Irreducible modular representations of
  {${\rm GL}_2$} of a local field}, Duke Math. J. \textbf{75} (1994), no.~2,
  261--292. \MR{1290194 (95g:22030)}

\bibitem[Her11a]{Her2011b}
Florian Herzig, \emph{The classification of irreducible admissible mod {$p$}
  representations of a {$p$}-adic {${\rm GL}_n$}}, Invent. Math. \textbf{186}
  (2011), no.~2, 373--434. \MR{2845621}

\bibitem[Her11b]{Her2011a}
\bysame, \emph{A {S}atake isomorphism in characteristic {$p$}}, Compos. Math.
  \textbf{147} (2011), no.~1, 263--283. \MR{2771132}

\bibitem[Le18]{Le2018}
Daniel Le, \emph{On some nonadmissible smooth irreducible representations for
  {$GL_2$}}, preprint, https://arxiv.org/abs/1809.10247, 2018.

\bibitem[Xu18]{X2018a}
Peng Xu, \emph{Hecke eigenvalues in {$p$}-modular representations of unramifed
  {${\rm U}(2,1)$}}, Preprint, 2018.

\end{thebibliography}

\documentclass[14pt,a4paper]{article}

\usepackage{mathrsfs}
\usepackage{amscd}
\usepackage{hyperref}
\hypersetup{colorlinks=true,citecolor=green,linkcolor=blue,urlcolor=blue}
\usepackage{tipa}
\usepackage{amssymb}
\usepackage{amsfonts}
\usepackage{stmaryrd}
\usepackage{amssymb}
\usepackage{titling}
\usepackage{CJK}  
\usepackage{indentfirst} 

\usepackage{latexsym}   
\usepackage{bm}         

\usepackage{pifont}
\usepackage{bm}
\usepackage{amsmath,amssymb,amsfonts}
\usepackage{indentfirst}

\usepackage{xcolor}
\usepackage{cases}
\usepackage{wasysym}
\usepackage{amsthm}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{remark}[theorem]{Remark}

\CJKtilde   

\begin{document}

\title{\large \textbf{Mod-$p$ maximal compact inductions do not have irreducible admissible subrepresentations}}
\date{}
\author{\textbf{Peng Xu}}
\maketitle

\begin{abstract}
Let $p$ be a prime number. We show in this short note that mod-$p$ maximal compact inductions of a $p$-adic split reductive group do not have irreducible admissible subrepresentations.
\end{abstract}

\section{Introduction}

Let $F$ be a non-archimedean local field of residue characteristic $p$, and $G$ be a $p$-adic split reductive group defined over $F$. Let $K$ be a hyperspecial maximal compact subgroup of $G$. Let $\sigma$ be an irreducible smooth $\overline{\mathbf{F}}_p$-representation of $K$. We show in this short note:

\begin{theorem}\label{main in intro}
The compactly induced representation $\textnormal{ind}^G _K \sigma$ does not have irreducible admissible subrepresentations.
\end{theorem}

\section{Proof of Theorem \ref{main in intro}}

Let $\sigma$ be an irreducible smooth $\overline{\mathbf{F}}_p$-representation of $K$.

\begin{proof}[Proof of Theorem \ref{main in intro}]

Assume $\pi$ is an irreducible admissible $\overline{\mathbf{F}}_p$-representation of $G$ contained in $\textnormal{ind}^G _K \sigma$, i.e., we are given a $G$-embedding
\begin{center}
$\iota: \pi \hookrightarrow \textnormal{ind}^G _K \sigma$.
\end{center}
Take an irreducible smooth $\overline{\mathbf{F}}_p$-representation $\sigma_1$ of $K$ contained in $\pi$. By Frobenius reciprocity, we get a non-zero $G$-map $\theta$ in the space $\text{Hom}_G (\textnormal{ind}^G _K \sigma_1, \pi)$, which will be denoted by $H(\sigma_1, \pi)$. Since $\pi$ is admissible, the space $H(\sigma_1, \pi)$ is finite dimensional. By composition, the space $H(\sigma_1, \pi)$ is a right module over the spherical Hecke algebra $\mathcal{H} (K, \sigma_1): =\text{End}_G (\textnormal{ind}^G _K \sigma_1)$. As $G$ is split, the algebra $\mathcal{H} (K, \sigma)$ is commutative (\cite[Corollary 1.3]{Her2011a}). Therefore we may replace $\theta$ by an eigenvector. That is to say, there is a character $\chi: \mathcal{H} (K, \sigma_1)\rightarrow \overline{\mathbf{F}}_p$ so that $\theta$ is annihilated by the kernel of $\chi$.

\smallskip

Now under our assumption, the composition $\iota\circ \theta$ is a non-zero map in $\text{Hom}_G (\textnormal{ind}^G _K \sigma_1, \textnormal{ind}^G _K \sigma)$. We take a non-zero $T$ in the kernel of $\chi$. We get
\begin{center}
$(\iota\circ \theta)\circ T =\iota \circ (\theta\circ T)=0$.
\end{center}
As $T$ and $\iota\circ \theta$ are both non-zero, we get a contradiction by the argument of \cite[Corollary 6.5]{Her2011b}.
\end{proof}

\begin{remark}
Note that we have assumed $\pi$ is admissible in the theorem. However, in certain cases such an assumption is not necessary, say $G= GL_2/ U(2, 1)$, as in both cases a weaker substitute, i.e., the existence of Hecke eigenvalues (\cite{B-L94}, \cite{X2018a}), is available. Note also that Daniel Le proved recently that there are non-admissible irreducible mod-$p$ smooth representations of $GL_2 (\mathbf{Q}_{p^3})$ (\cite{Le2018}).
\end{remark}

\begin{remark}
One interest of Theorem \ref{main in intro} is to compare it with the \textbf{complex} case: when $G$ is a classical group, a complex maximal compactly induced representation of $G$ might often happen to be irreducible.
\end{remark}

\bibliographystyle{amsalpha}
\bibliography{new}

\end{document}